OpenStudy (anonymous):

During a catered lunch, an average of 4 cups of tea are poured per minute. The lunch will last 2 hours. How many gallons of tea should the caterer bring if there are 16 cups in one gallon?

OpenStudy (anonymous):

4*60=240cups 240 \16=15

OpenStudy (anonymous):

15 gallons

OpenStudy (anonymous):

You figure it out c: \[4\times 120 = y\] \[y \div 16 = g\]

OpenStudy (anonymous):

Walter, it's for 2 hours.

OpenStudy (anonymous):

i got 30

OpenStudy (anonymous):

o sorry hold on

OpenStudy (anonymous):

The answer is 30 gallons of tea.

OpenStudy (anonymous):

4*120=480 480\16=30

OpenStudy (anonymous):

4 × 120 = 480 480 ÷ 16 = 30